OpenStudy (anonymous):

Rationalize the denominator of 5 over quantity of 9 plus square root of 5 quantity of 45 minus 5 square root of 5 over 76 quantity of 45 minus square root of 5 over 76 5 square root of 5 over quantity 9 square root of 5 end root plus 5 5 square root of 5 over 86

OpenStudy (anonymous):

5/ ( 9 + √5) Rationalize : 5 ( 9 - √5) / ( 9 + √5) ( 9 -√5) Can you continue?

OpenStudy (anonymous):

I really have no idea, I listened in class and everything but this one unit is what has me stumped, I understand if you dont want to give me the answer straight out answer but i really dont know what to do.

OpenStudy (anonymous):

( a + b) ( a - b) = a² - b² ( 9 + √5) ( 9 -√5) = ...?

OpenStudy (anonymous):

how do i add to the square root of 5? or even minus it..

OpenStudy (anonymous):

That's where you need to learn the formula to rationalize: ( a + b) ( a - b) = a² - b² a = 9, b = √5 ( 9 + √5) ( 9 -√5) = ...

OpenStudy (anonymous):

9 

OpenStudy (anonymous):

imean sqrt 14 and sqrt of 4 ?

OpenStudy (anonymous):

a = 9 --> a² = ...? b = √5 --->b² = ..?

OpenStudy (anonymous):

does a = 9^2 b = sqrt5^2 ?

OpenStudy (anonymous):

Yes, a² = 9² = ...?

OpenStudy (anonymous):

81

OpenStudy (anonymous):

b² = ...?

OpenStudy (anonymous):

sqrt 25

OpenStudy (anonymous):

b = √5 --->b² = ( √5 )² = 5

OpenStudy (anonymous):

Now a² - b² = ...?

OpenStudy (anonymous):

i dont know how to subtract that..

OpenStudy (anonymous):

a² - b² = 81 - ... =

OpenStudy (anonymous):

wait so 5?? 81 x 5?

OpenStudy (anonymous):

a² - b² = 81 - 5 =

OpenStudy (anonymous):

i mean minus... 76

OpenStudy (anonymous):

This is denominator: ( a + b) ( a - b) = a² - b² a = 9, b = √5 ( 9 + √5) ( 9 -√5) = 81- 5 = 76

OpenStudy (anonymous):

Numerator: 5 ( 9 - √5) = ...?

OpenStudy (anonymous):

32 1/2

OpenStudy (anonymous):

Distribute 5 into the bracket: 5 ( 9 - √5) = 45 - 5√5

OpenStudy (anonymous):

wouldnt that be like 33.something..

OpenStudy (anonymous):

Now compare my solution with the choices!

OpenStudy (anonymous):

wait the 5(9-sqrt5) one ?

OpenStudy (anonymous):

= ( 45 - 5√5 ) / 76

OpenStudy (anonymous):

answer a. thanks

OpenStudy (anonymous):

NO, it's not!

OpenStudy (anonymous):

what ?!

OpenStudy (anonymous):

I mean a choice is correct! About your asking if my solution as same as yours, so it's NOT!

OpenStudy (anonymous):

ohhk.